OpenStudy (anonymous):

Solve the quadratic equation x^2 - 4x + 57 = -5 by completing the square. A.x=1±i√50 B.x=2±√58 C.x=2±i√58 Dx=2±√62

OpenStudy (john_es):

\[ x^2 - 4x + 57 = -5\Rightarrow x^2 - 4x +4-4+ 57 = -5\Rightarrow\\ \Rightarrow (x - 2)^2-4 + 57 = -5\Rightarrow (x-2)^2=-58\]

OpenStudy (john_es):

\[x-2=\pm i\sqrt{58}\Rightarrow x=2\pm i\sqrt{58}\]

OpenStudy (shiraz14):

@John_ES is correct. Well Done!

OpenStudy (anonymous):

thank you !